1) Which of the following has a null value of ∆H◦f?

A)NO2
B)O2
C)CH4
D)H2O

2) A 200g piece of Al (Ce= 0.9J/g◦C) initially at 25 C◦ absorbs 20 kJ of heat. Calculate the final temperature of the metal?


3) A 2 moles methane (CH4) sample is placed in a bomb calorimeter with 50g of water and an initial temperature of 25 ◦C. A 4 moles methane (CH4) sample is placed in a bomb calorimeter with 50g of water and an initial temperature of 25 ◦C. After combustion of the gas, the water has a temperature of 48 ◦C. If the calorimeter constant was 17167 cal · C◦−1, calculate the combustion energy of CH4 in Kcal/mol?
1
Expert's answer
2020-04-24T13:36:10-0400

1)because entalpy fundamental simple is zero. O2

2)dT=Q*C*m

dt=111.1*C

Final temperature:136,1C

3)CH4+O2=CO2+2H2O

1 mole CH4=>2 mole H2O, 36 g water.

2 mole CH4=>4 mole H2O, 72 g

4 mole CH4=>8 mole H2O, 144g

Mass of water:50+72 =122or 50+144=194

The C_water=1cal/g*C

deltaT=23*C

Qwater=dT*C*m =2806 cal(or4462l

Qcalomiter=394703cal

Summ=397509 (or 399165)

Entalpy combustion reaction is Q/nmoles

1)198754,5=198,75Kcal/mol 2)99791,25cal=99,79Kcal/mol